# Gradient-based Sample Selection for Faster Bayesian Optimization

### v1

@@ A.2 Theorem 2:Analysis on Regret Bound:
$P\left(\left|f(\mathbf{x})-f([\mathbf{x}]_{n})\right|\leq brp\sqrt{\log\left(\frac{2pa}{\delta}\right)}/\tau_{n}\right)\geq 1-\frac{\delta}{2}\quad\forall\mathbf{x}\in\mathcal{X}_{n}.$
To cancel out the constants and keep the only dependence on stage $n$, we can set the discretization points $\tau_{n}=n^{2}brp\sqrt{\log\left(\frac{2pa}{\delta}\right)}$ along each dimension of the feature space, leading to:
$P\left(\left|f(\mathbf{x})-f([\mathbf{x}]_{n})\right|\leq\frac{1}{n^{2}}\right)\geq 1-\frac{\delta}{2}\quad\forall\mathbf{x}\in\mathcal{X}_{n},$
where the total number of discretization points becomes $|\mathcal{X}_{n}|=\left(n^{2}brp\sqrt{\log\left(\frac{2pa}{\delta}\right)}\right)^{p}$.
Now, using $\frac{\delta}{2}$ in lemma 4.3 and choosing $\mathbf{x}=[\mathbf{x}]_{n}\in\mathcal{X}_{n}$ gives:
$\left|f([\mathbf{x}^{*}]_{n})-\mu_{n-1}([\mathbf{x}^{*}]_{n})\right|=\left|f(\mathbf{x}^{*})-\mu_{n-1}([\mathbf{x}^{*}]_{n})+f([\mathbf{x}]_{n})-f(\mathbf{x}^{*})\right|\leq\left|f(\mathbf{x}^{*})-\mu_{n-1}([\mathbf{x}^{*}]_{n})\right|+\left|f([\mathbf{x}]_{n})-f(\mathbf{x}^{*})\right|,$
$\leq\tilde{\beta_{n}}^{1/2}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}.$
@@
$\tilde{\beta_{n}}=2\log\frac{|\mathcal{X}_{n}|\pi_{n}}{\delta/2}=2\log\frac{2\pi_{n}}{\delta}+2p\log\left(n^{2}brp\sqrt{\log\left(\frac{2pa}{\delta}\right)}\right),$
@@
For any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2p\log\left(n^{2}brp\sqrt{\log\left(\frac{4pa}{\delta}\right)}\right),i.e.,\beta_{n}=2\log\frac{4\pi_{n}}{\delta}+2p\log\left(n^{2}brp\sqrt{\log\left(\frac{4pa}{\delta}\right)}\right)-\frac{A}{\sigma_{\min}},$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1$, $\pi_{n}>0$. As $n\to\infty$, we have the following regret bound with probability $\geq 1-\delta$:
@@
$\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2p\log\left(n^{2}brp\sqrt{\log\frac{4pa}{\delta}}\right).$
@@
$\mu_{n-1}(\mathbf{x}_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\geq\mu_{n-1}([\mathbf{x}^{*}]_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n}).$
@@
$\mu_{n-1}([\mathbf{x}^{*}]_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}\geq f(\mathbf{x}^{*}).$
@@
$\mu_{n-1}(\mathbf{x}_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\geq(1-d)\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+f(\mathbf{x}^{*})-\frac{1}{n^{2}}.$
@@
$r_{t}=f(\mathbf{x}^{*})-f(\mathbf{x}_{n})\leq\mu_{n-1}(\mathbf{x}_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})+\frac{1}{n^{2}}+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n}),$
$\leq(d+1)\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})+(d-1)\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}.$
@@
$r_{n}\leq 2\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n}).$
@@
The mutual information gain for a total of $N$ stages can be expressed as follows:
$I(y_{1:N};f_{1:N})=\frac{1}{2}\sum_{n=1}^{N}\log\left(1+\sigma^{-2}\sigma_{n-1}^{2}(\mathbf{x}_{n})\right)$
@@
Recall that $I(y_{1:N};f_{1:N})=H(y_{1:N})-\frac{1}{2}\log|2\pi e\sigma^{2}\mathbf{I}|$. Using the chain rule of conditional entropy gives:
$H(y_{1:N})=H(y_{1:N-1})+H(y_{1:N}|y_{1:N-1})$
$=H(y_{1:N-1})+\frac{1}{2}\log(2\pi e(\sigma^{2}+\sigma_{N-1}^{2}(\mathbf{x}_{N})))$
@@
$I(y_{1:N};f_{1:N})=H(y_{1:N-1})+\frac{1}{2}\log(2\pi e(\sigma^{2}+\sigma_{N-1}^{2}(\mathbf{x}_{N})))-\frac{1}{2}\log|2\pi e\sigma^{2}\mathbf{I}|$
$=H(y_{1:N-1})+\frac{1}{2}\log\left(1+\sigma^{-2}\sigma_{N-1}^{2}(\mathbf{x}_{N})\right)$
Note that $\mathbf{x}_{1},\ldots,\mathbf{x}_{N}$ are deterministic given the outcome observations $y_{1:N-1}$, and the conditional variance term $\sigma_{N-1}^{2}(\mathbf{x}_{N})$ does not depend on the realization of $y_{1:N-1}$ due to the conditioning property of the GP. The result thus follows by induction. ∎
Now we provide proof for the main theorem on the the regret bound. We use $\mathcal{O}^{*}$, a variant of the $\mathcal{O}$ notation to suppress the log factors.
@@
Based on 4.5, we have $r_{n}^{2}\leq 2\tilde{\beta_{n}}\sigma_{n-1}^{2}(\mathbf{x}_{n})$ with probability $\geq 1-\delta$ as $n\to\infty$. Since $\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2p\log\left(n^{2}brp\sqrt{\log\frac{4pa}{\delta}}\right)$ and is nondecreasing in $n$, we can upper bound it by the final stage $\tilde{\beta_{N}}$:
$2\tilde{\beta_{n}}\sigma_{n-1}^{2}(\mathbf{x}_{n})\leq 2\tilde{\beta_{N}}\sigma^{2}\frac{\sigma_{n-1}^{2}(\mathbf{x}_{n})}{\log(1+\sigma^{-2})}=2\tilde{\beta_{N}}\sigma^{2}C^{\prime\prime}\log(1+\sigma^{-2}),$
where $C^{\prime\prime}=\frac{\sigma^{-2}}{\log(1+\sigma^{-2})}$.
@@
$R_{N}^{2}\leq N\sum_{n=1}^{N}r_{n}^{2}\leq N\sum_{n=1}^{N}2\beta_{N}\sigma^{2}C^{\prime\prime}\log(1+\sigma^{-2}),$
$=4N\beta_{N}\sigma^{2}C^{\prime\prime}I(y_{1:N};f_{1:N}),$
$=C_{1}N\beta_{N}I(y_{1:N};f_{1:N}),$
$\leq C_{1}N\beta_{N}\gamma_{N},$
where $C=8\sigma^{2}C^{\prime\prime}$ and $\gamma_{N}=\max I(y_{1:N};f_{1:N})$ is the maximum information gain after $N$ steps of sampling. Thus,
$P\left(R_{N}\leq\sqrt{CN\beta_{N}\gamma_{N}}\right)\geq 1-\delta.$
Note that our main theorem’s form is quite similar with {Srinivas et al., 2010}, although our stage-wise constant $\beta_{n,d}$ is different and includes a distance term.
@@ A.3 Theorem 3: Two-Phase Regret
Let $N$ be the total number of rounds. Suppose the first $M_{\text{initial}}$ rounds use the full GP-UCB strategy (i.e., no sparse approximation), while rounds $t=1$ to $t=N$ employ a GP-UCB strategy. The cumulative regret denote by $R_{N}\;=\;\sum_{t=1}^{N}\bigl{(}f(\mathbf{x}^{*})-f(\mathbf{x}_{t})\bigr{)}$ , where $\mathbf{x}^{*}$ is an optimal point and $\mathbf{x}_{t}$ is the decision made at time $t$. Decompose the $N$ rounds into two segments:
$R_{N}\;=\;\underbrace{\sum_{t=1}^{M}\bigl{(}f(\mathbf{x}^{*})-f(\mathbf{x}_{t})\bigr{)}}_{\displaystyle R_{M}^{(\text{full})}}\;+\;\underbrace{\sum_{t=M+1}^{T}\bigl{(}f(\mathbf{x}^{*})-f(\mathbf{x}_{t})\bigr{)}}_{\displaystyle R_{N-M}^{(\text{selected})}}.$
1. Regret Bound in the First $M$ Rounds
During the initial $M$ rounds, the strategy relies on the standard GP-UCB. By the well-known GP-UCB regret bounds (Srinivas et al., 2009), there exists a constant $C_{1}$, pick $\delta\in(0,1)$, and define $\beta_{t}=2\log\left(t^{2}2\pi^{2}/(3\delta)\right)+2p\log\left(t^{2}brp\sqrt{\log(4da/\delta)}\right)$, we have,
$\Pr\left\{R_{M}^{(\text{full})}\leq\sqrt{CM\beta_{M}\gamma_{M}+2}\quad\forall M\geq 1\right\}\geq 1-\delta.$
2. Regret Bound from Round $M+1$ to $N$
Starting from iteration $t=M+1$, the regret analysis switches to a sparse GP-UCB. Let $R_{N-M}^{(\text{selected})}$ denote the regret incurred in these final $N-M$ rounds.
Choose $\beta_{n}=2\log\frac{4\pi_{n}}{\delta}+2p\log\left(n^{2}brp\sqrt{\log\left(\frac{4pa}{\delta}\right)}\right)-\frac{A}{\sigma_{\min}},$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\quad\pi_{n}>0$. As $n\to\infty$, we obtain a regret bound of $\mathcal{O}^{*}(\sqrt{p(N-M)\gamma_{(N-M)}})$. Specifically, with $C_{1}=\frac{8}{\log(1+\sigma^{-2})},$ we have:
$\Pr\left(R_{N-M}\leq\sqrt{C(N-M)\beta_{(N-M)}\gamma_{(N-M)}}\quad\forall{N-M}\geq 1\right)\geq 1-\delta.$

### v2

@@ A.2 Theorem 2:Analysis on Regret Bound:
$P\left(\left|f(\mathbf{x})-f([\mathbf{x}]_{n})\right|\leq brd\sqrt{\log\left(\frac{2da}{\delta}\right)}/\tau_{n}\right)\geq 1-\frac{\delta}{2}\quad\forall\mathbf{x}\in\mathcal{X}_{n}.$ (39)
To cancel out the constants and keep the only dependence on stage $n$, we can set the discretization points $\tau_{n}=n^{2}brd\sqrt{\log\left(\frac{2da}{\delta}\right)}$ along each dimension of the feature space, leading to:
$P\left(\left|f(\mathbf{x})-f([\mathbf{x}]_{n})\right|\leq\frac{1}{n^{2}}\right)\geq 1-\frac{\delta}{2}\quad\forall\mathbf{x}\in\mathcal{X}_{n},$ (40)
where the total number of discretization points becomes $|\mathcal{X}_{n}|=\left(n^{2}brd\sqrt{\log\left(\frac{2da}{\delta}\right)}\right)^{d}$. Now, using $\frac{\delta}{2}$ in lemma 4.3 and choosing $\mathbf{x}=[\mathbf{x}]_{n}\in\mathcal{X}_{n}$ gives:
$\begin{split}\bigl{|}f([\mathbf{x}^{*}]_{n})-\mu_{n-1}([\mathbf{x}^{*}]_{n})\bigr{|}&=\bigl{|}f(\mathbf{x}^{*})-\mu_{n-1}([\mathbf{x}^{*}]_{n})+f([\mathbf{x}]_{n})-f(\mathbf{x}^{*})\bigr{|}\\[6.0pt] &\leq\bigl{|}f(\mathbf{x}^{*})-\mu_{n-1}([\mathbf{x}^{*}]_{n})\bigr{|}+\bigl{|}f([\mathbf{x}]_{n})-f(\mathbf{x}^{*})\bigr{|}\\[6.0pt] &\leq\tilde{\beta}_{n}^{1/2}\,\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}.\end{split}$ (41)
@@
$\tilde{\beta_{n}}=2\log\frac{|\mathcal{X}_{n}|\pi_{n}}{\delta/2}=2\log\frac{2\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{2da}{\delta}\right)}\right),$ (42)
@@
For any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{4da}{\delta}\right)}\right),i.e.,$ $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,\Bigl{[}2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{4da}{\delta}\right)}\right)\Bigr{]}\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1$, $\pi_{n}>0$. As $n\to\infty$, we have the following regret bound with probability $\geq 1-\delta$:
@@
$\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\frac{4da}{\delta}}\right).$ (43)
@@
$\mu_{n-1}(\mathbf{x}_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\geq\mu_{n-1}([\mathbf{x}^{*}]_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n}).$ (44)
@@
$\mu_{n-1}([\mathbf{x}^{*}]_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}\geq f(\mathbf{x}^{*}).$ (45)
@@
$\mu_{n-1}(\mathbf{x}_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\geq(1-d)\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+f(\mathbf{x}^{*})-\frac{1}{n^{2}}.$ (46)
@@
$\begin{split}r_{t}=f(\mathbf{x}^{*})-f(\mathbf{x}_{n})&\leq\mu_{n-1}(\mathbf{x}_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\,\sigma_{n-1}(\mathbf{x}_{n})+\frac{1}{n^{2}}+\tilde{\beta_{n}}^{\frac{1}{2}}\,\sigma_{n-1}([\mathbf{x}^{*}]_{n}),\\ &\leq(d+1)\,\tilde{\beta_{n}}^{\frac{1}{2}}\,\sigma_{n-1}(\mathbf{x}_{n})+(d-1)\,\tilde{\beta_{n}}^{\frac{1}{2}}\,\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}.\end{split}$ (47)
@@
$r_{n}\leq 2\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n}).$ (48)
@@
The mutual information gain for a total of ${T_{N}}$ stages can be expressed as follows:
$I(y_{1:{T_{N}}};f_{1:{T_{N}}})=\frac{1}{2}\sum_{n=1}^{T_{N}}\log\left(1+\sigma^{-2}\sigma_{n-1}^{2}(\mathbf{x}_{n})\right)$ (49)
@@
Recall that $I(y_{1:{T_{N}}};f_{1:{T_{N}}})=H(y_{1:{T_{N}}})-\frac{1}{2}\log|2\pi e\sigma^{2}\mathbf{I}|$. Using the chain rule of conditional entropy gives:
$\begin{split}H\bigl{(}y_{1:T_{N}}\bigr{)}&=H\bigl{(}y_{1:T_{N-1}}\bigr{)}+H\bigl{(}y_{1:T_{N}}\mid y_{1:T_{N-1}}\bigr{)}\\[6.0pt] &=H\bigl{(}y_{1:T_{N-1}}\bigr{)}+\frac{1}{2}\,\log\!\bigl{(}2\pi e\,\bigl{(}\sigma^{2}+\sigma_{T_{N-1}}^{2}(\mathbf{x}_{T_{N}})\bigr{)}\bigr{)}\,.\end{split}$ (50)
@@
$\begin{split}I\bigl{(}y_{1:T_{N}};\,f_{1:T_{N}}\bigr{)}&=H\bigl{(}y_{1:T_{N-1}}\bigr{)}+\frac{1}{2}\log\bigl{(}2\pi e\,[\sigma^{2}+\sigma_{T_{N-1}}^{2}(\mathbf{x}_{T_{N}})]\bigr{)}-\frac{1}{2}\log\bigl{|}2\pi e\,\sigma^{2}\mathbf{I}\bigr{|}\\[6.0pt] &=H\bigl{(}y_{1:T_{N-1}}\bigr{)}+\frac{1}{2}\log\!\Bigl{(}1+\sigma^{-2}\,\sigma_{T_{N-1}}^{2}(\mathbf{x}_{T_{N}})\Bigr{)}\end{split}$ (51)
Note that $\mathbf{x}_{1},\ldots,\mathbf{x}_{T_{N}}$ are deterministic given the outcome observations $y_{1:T_{N-1}}$, and the conditional variance term $\sigma_{T_{N-1}}^{2}(\mathbf{x}_{T_{N}})$ does not depend on the realization of $y_{1:T_{N-1}}$ due to the conditioning property of the GP. The result thus follows by induction. ∎
Now we provide proof for the main theorem on the regret bound. We use $\mathcal{O}^{*}$, a variant of the $\mathcal{O}$ notation to suppress the log factors.
@@
Based on 4.5, we have $r_{n}^{2}\leq 2\tilde{\beta_{n}}\sigma_{n-1}^{2}(\mathbf{x}_{n})$ with probability $\geq 1-\delta$ as $n\to\infty$. Since $\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\frac{4da}{\delta}}\right)$ and is nondecreasing in $n$, we can upper bound it by the final stage $\tilde{\beta_{N}}$:
$2\tilde{\beta_{n}}\sigma_{n-1}^{2}(\mathbf{x}_{n})\leq 2\tilde{\beta_{N}}\sigma^{2}\frac{\sigma_{n-1}^{2}(\mathbf{x}_{n})}{\log(1+\sigma^{-2})}=(1/4)\tilde{\beta_{N}}C_{1}\log(1+\sigma^{-2}).$ (52)
@@
$\begin{split}R_{T_{N}}^{2}&\leq T_{N}\sum_{n=1}^{T_{N}}r_{n}^{2}\;\leq\;T_{N}\sum_{n=1}^{T_{N}}\frac{1}{4}\,\beta_{T_{N}}\,\log\bigl{(}1+\sigma^{-2}\bigr{)}\\[6.0pt] &=C_{1}\,T_{N}\,\beta_{T_{N}}\,I\bigl{(}y_{1:T_{N}};\,f_{1:T_{N}}\bigr{)}\\[6.0pt] &\leq C_{1}\,T_{N}\,\beta_{T_{N}}\,\gamma_{T_{N}}\,.\end{split}$ (53)
where $\gamma_{T_{N}}=\max I(y_{1:{T_{N}}};f_{1:{T_{N}}})$ is the maximum information gain after ${T_{N}}$ steps of sampling. Thus,
$P\left(R_{T_{N}}\leq\sqrt{C_{1}{T_{N}}\beta_{{T_{N}}}\gamma_{T_{N}}}\right)\geq 1-\delta.$ (54)
Note that our main theorem’s form is quite similar to {Srinivas et al., 2010}, although our stage-wise constant $\beta_{n,d}$ is different and includes a distance term.
@@ A.3 Theorem 3: Two-Phase Regret
Let $N$ be the total number of rounds. Suppose the first ${T_{M}}$ rounds use the full GP-UCB strategy, while rounds $t={T_{M+1}}$ to $t={T_{N}}$ employ a GSSBO strategy. The cumulative regret is denoted by $R_{T_{N}}\;=\;\sum_{t=1}^{T_{N}}\bigl{(}f(\mathbf{x}^{*})-f(\mathbf{x}_{t})\bigr{)}$ , where $\mathbf{x}^{*}$ is an optimal point and $\mathbf{x}_{t}$ is the decision made at time $t$. Decompose the $N$ rounds into two segments:
$R_{T_{N}}\;=\;\underbrace{\sum_{t=1}^{{T_{M}}}\bigl{(}f(\mathbf{x}^{*})-f(\mathbf{x}_{t})\bigr{)}}_{\displaystyle R_{{T_{M}}}^{(\text{full})}}\;+\;\underbrace{\sum_{t={T_{M}}+1}^{{T_{N}}}\bigl{(}f(\mathbf{x}^{*})-f(\mathbf{x}_{t})\bigr{)}}_{\displaystyle R_{{T_{N}}-{T_{M}}}^{(\text{selected})}}.$ (55)
1. Regret Bound in the First ${T_{M}}$ Rounds
During the initial $M$ rounds, the strategy relies on the standard GP-UCB. By the well-known GP-UCB regret bounds [46], there exists a constant $C_{1}=\frac{8}{\log(1+\sigma^{-2})}$, pick $\delta\in(0,1)$, and define $\beta_{n}=2\log\left(n^{2}2\pi^{2}/(3\delta)\right)+2d\log\left(n^{2}brd\sqrt{\log(4da/\delta)}\right)$, we have,
$\Pr\left\{R_{M}^{(\text{full})}\leq\sqrt{C_{1}M\beta_{M}\gamma_{M}}+2\quad\forall M\geq 1\right\}\geq 1-\delta.$ (56)
2. Regret Bound from Round ${T_{M}}+1$ to ${T_{N}}$
Starting from iteration $t={T_{M}}+1$, the regret analysis switches to a sparse GP-UCB. Let $R_{{T_{N}}-{T_{M}}}^{(\text{selected})}$ denote the regret incurred in these final ${T_{N}}-{T_{M}}$ rounds. Choose $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,\Bigl{[}2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{4da}{\delta}\right)}\right)\Bigr{]}\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\quad\pi_{n}>0$. As $n\to\infty$, we obtain a regret bound of $\mathcal{O}^{*}(\sqrt{d({T_{N}}-{T_{M}})\gamma_{({T_{N}}-{T_{M}})}})$. Specifically, we have:
$\Pr\left(R_{{T_{N}}-{T_{M}}}\leq\sqrt{C_{1}({T_{N}}-{T_{M}})\beta_{({T_{N}}-{T_{M}})}\gamma_{({T_{N}}-{T_{M}})}}\quad\forall{({T_{N}}-{T_{M}})}\geq 1\right)\geq 1-\delta.$ (57)